OpenStudy (moazzam07):

At Sami's Shoe Warehouse, it takes fraction 4 over 5 of a day to complete fraction 1 over 10 of an order of sneakers. At this rate, how long will it take to complete the entire order of sneakers? 6 days 8 days Mixed number 11 and 1 over 4 days Mixed number 12 and 1 over 2 days

OpenStudy (anonymous):

Set up the two fractions together like this

OpenStudy (anonymous):

Sorry, my equation tool isn't working

OpenStudy (anonymous):

Set up 4/5 = 1/10

OpenStudy (moazzam07):

ok

OpenStudy (anonymous):

The left is the amount of time it takes to do the work, the right is the amount of work done right? So you want your 1/10 to equal 10/10 or 1

OpenStudy (moazzam07):

i dont get it

OpenStudy (moazzam07):

oh ok

OpenStudy (moazzam07):

i got it

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 } = \frac{ 1 }{ 10 }\]

OpenStudy (moazzam07):

so what should i do to turn 1/10 into 10/10

OpenStudy (anonymous):

What can you do to 1/10 to make it equal to 1?

OpenStudy (anonymous):

Well if 1/10 means one part of ten, you would want to multiply it by ten

OpenStudy (moazzam07):

yes

OpenStudy (anonymous):

And since we have 4/5 and 1/10 as a part of the same equation, we have to do the same thing to the left as we do to the right

OpenStudy (moazzam07):

ok

OpenStudy (moazzam07):

so multiply both sides by 10?

OpenStudy (anonymous):

So\[\frac{ 1 }{ 10 } \times \frac{ 10 }{ 1 } = \frac{ 4 }{ 5 } \times \frac{ 10 }{ 1 }\]

OpenStudy (moazzam07):

1 = 40/5

OpenStudy (anonymous):

Yes, which should reduce to 8

OpenStudy (moazzam07):

1=8

OpenStudy (anonymous):

And we remember the one should now equal one load of sneakers equals eight days of work

OpenStudy (moazzam07):

Thank you, i have 24 more questions can you help?

OpenStudy (anonymous):

I can offer advice, but I won't solve them all for you. Find one of the harder ones

OpenStudy (moazzam07):

ok

OpenStudy (moazzam07):

got it

OpenStudy (moazzam07):

What is the first step needed to solve 3 over 4 the whole multiplied by x minus 3 equal negative 18? Add 3 to both sides Subtract 18 from both sides Multiply both sides by 3 Divide both sides by 4

OpenStudy (moazzam07):

i think its add 3 to both sides

OpenStudy (moazzam07):

\[\frac{ 3 }{ 4 } x - 3 - 18\]

OpenStudy (moazzam07):

i mean 3 = -18

OpenStudy (anonymous):

Then yes, it should be add 3 to both sides

OpenStudy (moazzam07):

yay ok

OpenStudy (moazzam07):

then i got Which fraction is equivalent to negative 2 over 3? 2 over negative 3 negative 3 over 2 negative 2 over negative 3 2 over 3

OpenStudy (anonymous):

Negative signs on fractions can be pulled out. Think of it as a division problem instead of a fraction -2/3 is going to return a negative number, right?

OpenStudy (anonymous):

Two negatives and two positives both return a positive number

OpenStudy (moazzam07):

2 over 3

OpenStudy (anonymous):

Nope, 2 divided by three would be positive

OpenStudy (anonymous):

If there is one negative in the equation, it will be negative, so it should be 2 over negative 3